OpenStudy (anonymous):

A 5.0-kg bucket of water is raised from a well by a rope. If the upward acceleration of the bucket is 3.0 m/s2, find the force exerted by the rope on the bucket.

OpenStudy (anonymous):

newtons second law tells you sum of forces = ma the forces acting on the bucket are the force of the rope up and gravity down, these combine to create an upward accel of 3 m/s/s so we have: sum of forces = F-mg = ma F-5g=5x3m/s/s F=5x9.8+5x3 = 49+15=64N

OpenStudy (anonymous):

@undeadknight26 leggo lol